OpenStudy (bossimbacon):

Which is equal to 8^-10/8^5? A. 8^-50 B. 8^-15 C. 8^-5 D. 8^-2 @gorv

OpenStudy (anonymous):

\[\frac{ x^n }{ x^m } = x ^{n-m}\]

OpenStudy (bossimbacon):

so b

OpenStudy (anonymous):

\[\frac{ 8^{-10} }{ 8^5 } = 8^{-10-5} = 8^{-15}\]

OpenStudy (anonymous):

Yup.